# Trainable Dynamic Mask Sparse Attention

### v1

@@ 3.2 Multi-Head Attention
QKV Projection. In the Transformer, we first convert the input into query, key, and value representations. For the hidden state $h_{t}\in\mathbb{R}^{d_{model}}$ of the $t$-th token, we perform linear projections through the weight matrices $W^{Q}$, $W^{K}$, and $W^{V}$ to obtain $q_{t}$, $k_{t}$, and $v_{t}$ respectively, as shown in Equation 1. These projections transform the input representation into separate subspaces for each attention head, allowing each head to focus on distinct aspects of the input. The weight matrices shape the projection into $n_{h}$ attention heads, each with a dimension of $d_{h}$, where $d_{model}=n_{h}\times d_{h}$ is typically held.
@@
Key-Value Concatenation. In autoregressive generation, we need to cache the key-value pairs of historical tokens to avoid redundant computations. As shown in Equation 2, we concatenate the cached past token key matrix and value matrix with the current token’s key-value representations to form the complete key matrix $k$ and value matrix $v$. By maintaining and updating this cache, we construct a complete context window that spans all tokens from position 1 to the current position $t$, enabling the model to access and utilize the full sequence history.
$k$ $=\operatorname{concat}([k_{1},\ldots,k_{t}])\quad where\quad k\in\mathbb{R}^{n_{h}\times t\times d_{h}}$ (2)
$v$ $=\operatorname{concat}([v_{1},\ldots,v_{t}])\quad where\quad v\in\mathbb{R}^{n_{h}\times t\times d_{h}}$
@@ 4 Method
Figure 3: Dynamic Mask Attention Structure. It demonstrates the mask structure and weight structure of Dynamic Mask Attention in the multi-head case. Unlike the same and redundant mask and weight structures in Self-Attention and State-Space, the mask structure of Dynamic Mask Attention is dynamically adjusted through content awareness, where each head’s mask can be different. This allows Dynamic Mask Attention to achieve different attention weight distributions in each head, enabling the model to maximize the utilization of each subspace in multi-head attention and focus on different tokens in each head.
@@
As discussed in Section 2, existing attention mechanisms face numerous challenges when processing long sequences, including high computational complexity, large memory requirements, limitations of static masks, and inadequate handling of positional information. Dynamic Mask Attention cleverly leverages these sparsity patterns in language modeling, aiming to combine the advantages of State-Space and Self-Attention algorithms through two components: content-aware dynamic sparse masks and position-aware sparse attention weight computation, as shown in Figure 1. The former utilizes value representations to dynamically generate masks that determine which historical tokens each attention head should focus on, while the latter performs efficient sparse attention computations based on these masks. As shown in Figure 3, unlike traditional methods, DMA generates unique mask structures for each attention head, enabling the model to efficiently capture diverse content patterns across different subspaces. We provide a sample PyTorch implementation in Listing 8 for reference.
Content-Aware Dynamic Sparse Mask.
The content-aware dynamic sparse mask is the core innovation of DMA, which determines the historical information relevant to the current query by analyzing the content features of value representations. This approach differs from static masks based on predefined rules, as it can dynamically adapt to the varying needs of different contexts.
We first compute the dynamic attention weights $\delta$, as shown in Equation 3. Unlike traditional attention mechanisms that rely solely on query-key similarity, DMA introduces a content-aware dynamic weight generation mechanism based on value representations, which is a learnable way to determine which historical information should be retained.
$\delta$ $=\exp(\tau(v\Delta)\times A)\quad where\quad\Delta\in\mathbb{R}^{n_{h}\times d_{h}\times n_{h}}\quad A\in\mathbb{R}^{n_{h}}\quad\delta\in\mathbb{R}^{n_{h}\times t}$ (3)
Specifically, through the value matrix $v$, sampling weight matrix $\Delta$, and gating parameter $A$, we generate a set of adaptive weight factors. The $\Delta$ parameter acts similarly to a forget gate, controlling the degree of attention to the current input: larger $\Delta$ values reset the state and focus on the current input, while smaller $\Delta$ values maintain the state and ignore the current input. The zero-order hold method ensures stability of weights across sequence lengths, while the gating parameter $A$ provides fine-grained selective control. The non-negative function $\tau(\cdot)$ ensures that the final weights emphasize rather than suppress attention signals.
The key advantage of this method is that it achieves content-aware selectivity: for important input tokens, the model generates larger dynamic weights to focus on them; for less important tokens, it generates smaller weights to selectively ignore them. This mechanism allows the model to adaptively adjust the attention distribution based on the content of the input, achieving a similar effect to assigning different weights to different tokens based on their importance in the Transformer attention mechanism, but through a more efficient sparsification approach. At the same time, this design reserves a path for positional encoding, supporting a Self-Attention mechanism that is more adaptable to dynamic sparsity.
Then, we combine the attention bias with a causal mask to enforce autoregressive properties and perform sparse selection, as shown in Equation 4. Specifically, the dynamic mask $m_{t}$ is generated by adding the attention bias $\delta$ to the causal mask $m_{c}$, followed by applying a top-k operation and a sparsification function $f(\cdot)$. The top-k operation examines the weight vector $\delta_{n_{h},\cdot}+m_{\cdot}^{c}$ for each attention head $n_{h}$ when the sequence length $t$ exceeds $w$, retaining the top $w$ values in this vector and assigning them to the corresponding positions in $m^{d}_{n_{h},\cdot}$, while setting other positions to $\infty$- ∞, thus achieving sparsification. The combination of these two masks $m_{t}$ creates a two-dimensional tensor of shape $\mathbb{R}^{n_{h}\times t}$, where each attention head $n_{h}$ has a specific attention pattern.
$m_{t}$ $=f(top_{w}(\delta+m_{t}^{c}))$ (4)
@@
$where\quad m_{t}^{c}\in\mathbb{R}^{t}\quad m_{t}\in\mathbb{R}^{n_{h}\times t}$
Position-Aware Sparse Attention Weights.
The position-aware sparse attention weights computation is the second core component of DMA, which achieves efficient sparse attention weights computation by utilizing dynamically generated masks. This method allows for different attention weight distributions in each attention head, enabling the model to maximize the utilization of each subspace in multi-head attention and focus on different tokens in each head.
@@
The entire computation flow is shown in Equation LABEL:eq:attention_output. For each attention head $n_{h}$, we compute the scaled dot product $\frac{q_{t}k^{\top}}{\sqrt{d_{h}}}$ between the query and key, and then perform an element-wise $\circ$ operation with the combined mask $m_{t}$ constructed earlier. The scaling factor $\sqrt{d_{h}}$ is crucial here as it prevents the magnitude of the dot product from becoming too large, avoiding pushing the softmax function into a gradient-saturated region. After applying the mask, the softmax function normalizes the result, generating the attention weights $a_{n_{h},j}$. Notably, when the mask value $m_{n_{h},j}=-\infty$, the corresponding attention weight $a_{n_{h},j}\approx 0$, so we effectively skip the masked positions and fill that position with 0, allowing the model to focus only on the relevant context that is not masked. The attention weights for each head are then multiplied by the value vectors and summed to produce context vectors $o_{t}\in\mathbb{R}^{n_{h}\times d_{h}}$, where each row captures different contextual patterns and dependencies. The multi-head mechanism, combined with dynamic attention masks, enables the model to focus on multiple patterns in the sequence simultaneously. These context vectors are concatenated and projected back to the model dimension through an output projection matrix $W^{OD}\in\mathbb{R}^{n_{h}\times d_{h}\times d_{model}}$, resulting in the final attention output $h^{\prime}_{t}\in\mathbb{R}^{d_{model}}$. This output integrates information from all attention heads, forming a rich contextual representation that effectively captures different distance dependencies in the sequence history while maintaining computational efficiency.
@@
$o_{t}$ $=\operatorname{softmax}(\frac{q_{t}k^{\top}}{\sqrt{d_{h}}}\circ m_{t})v$ (5)

### v6

@@ 3.2 Multi-Head Attention
QKV Projection. In the Transformer architecture [vaswani2017attention], the input is first transformed into $Q$, $K$, $V$. For the hidden state of the $t$-th token in a sequence of length $n$, denoted as $h_{t}\in\mathbb{R}^{d_{model}}$, the linear projections are performed using weight matrices $W^{Q}$, $W^{K}$, and $W^{V}$ to obtain $q_{t}$, $k_{t}$, and $v_{t}$, respectively, as shown in Equation 1. These projections map the input representation into distinct subspaces for each of the $n_{h}$ attention heads, allowing each head to focus on different aspects of the input. The weight matrices shape the projections to have a dimension of $d_{h}$ per head.
@@
Key-Value Concatenation. During autoregressive generation, the key-value pairs of historical tokens are cached to prevent redundant computations. As shown in Equation 2, the cached key and value matrices from past tokens are concatenated with the key-value representations of the current token to form the complete key matrix $K$ and value matrix $V$. By maintaining and updating this cache, a complete context window spanning all tokens from position 1 to the current position $t$ is constructed, enabling the model to access and utilize the full sequence history.
$k$ $=\operatorname{concat}([k_{1},\ldots,k_{t}])\quad where\quad k\in\mathbb{R}^{n_{h}\times n\times d_{h}}$ (2)
$v$ $=\operatorname{concat}([v_{1},\ldots,v_{t}])\quad where\quad v\in\mathbb{R}^{n_{h}\times n\times d_{h}}$
@@ 4 Method
Figure 4: Dynamic Mask Attention Structure. It demonstrates the mask structure and weight structure of Dynamic Mask Attention in the multi-head case. Unlike the same and redundant mask and weight structures in Self-Attention and State-Space, the mask structure of DMA is dynamically adjusted through content awareness, where each head’s mask can be different. This allows DMA to achieve different attention weight distributions in each head, enabling the model to maximize the utilization of each subspace in multi-head attention and focus on different tokens in each head.
@@
As discussed in Section 2, existing attention mechanisms are confronted with numerous challenges when processing long sequences, including high computational complexity, substantial memory requirements, the rigidity of static masks, and the presence of non-differentiable components that impede end-to-end training. To address these issues, we introduce Dynamic Mask Attention, a mechanism that strategically leverages the inherent sparsity patterns of language modeling. As illustrated in Figure 2, DMA is composed of two core components: content-aware dynamic sparse masking and position-aware sparse attention weight computation. The former utilizes value representations to dynamically generate masks that determine which historical tokens each attention head should attend to, while the latter performs efficient sparse attention computations guided by these masks. This dual-component design allows DMA to maintain focus on critical information while adapting to varying contextual dependencies. Furthermore, as shown in Figure 4, DMA generates a unique mask structure for each attention head, enabling the model to capture diverse content patterns across different representational subspaces. A sample PyTorch implementation is provided in Listing LABEL:lst:dynamic_mask_attn_code for reference.
4.1 Generate Content-Aware Dynamic Mask
The content-aware dynamic mask constitutes the central innovation of DMA. It operates by analyzing the content features embedded within the value representations to determine which historical information is relevant to the current query. For each attention head at the current time step, this mechanism generates a unique dynamic mask that directs the subsequent attention weight computation to focus exclusively on the most critical key positions.
To sample the value vector representations, we introduce a sampling weight matrix $\Delta\in\mathbb{R}^{n_{h}\times d_{h}\times n_{h}}$, a per-head gating coefficient $A\in\mathbb{R}^{n_{h}}$, and a non-negative activation function $\tau(\cdot)$. As detailed in Equation (3), the process begins with a tensor contraction, $v\cdot\Delta$, which projects each token’s $d_{h}$-dimensional value vector into a scalar representation, serving as an initial estimate of its importance. Subsequently, the activation function $\tau(\cdot)$ ensures these scores are non-negative, preventing signal cancellation. The gating coefficient $A$ then scales the importance scores for each head, allowing the model to learn varying degrees of sparsity across heads, where $A$ can be made query-dependent, enabling the sampling scores to adapt based on the input. Finally, an exponential function amplifies the differences between scores and maps them to a positive value space, which facilitates the learning of gating effects and yields the final scores, $\delta\in\mathbb{R}^{n_{h}\times n}$.
$\delta$ $=\exp(\tau(v\cdot\Delta)\times A)\quad where\quad\Delta\in\mathbb{R}^{n_{h}\times d_{h}\times n_{h}}\quad A\in\mathbb{R}^{n_{h}}\quad\delta\in\mathbb{R}^{n_{h}\times n}$ (3)
Subsequently, as defined in Equation (4), a sparsification function $f(\cdot)$ is applied. This function identifies whether each score $\delta_{n_{h},j}$ ranks within the $top_{w}$ for its respective head. Scores within the $top_{w}$ are retained to preserve the gradient flow, while all other scores are set to -$\infty$, effectively nullifying their contribution in the subsequent softmax operation. For causal language modeling, this function can incorporate a causal mask via broadcasting to avoid additional memory overhead. This process yields the final dynamic mask, $m_{t}\in\mathbb{R}^{n_{h}\times n}$.
$m_{t}$ $=f(\delta)\quad where\quad m_{t}\in\mathbb{R}^{n_{h}\times n}$ (4)
@@ 4.1 Generate Content-Aware Dynamic Mask
This approach offers three significant advantages. First, by sampling importance scores from value representations, the model can more accurately focus on semantically critical tokens, regardless of their distance. This mitigates the risk of overlooking important long-range dependencies, a common issue with methods relying solely on positional patterns. Second, the combination of the gating coefficient $A$ and independent $top_{w}$ selection allows different attention heads to specialize in distinct functions, such as local, long-range, and global, thereby improving the breadth of representational coverage. Third, the sparse selection mechanism is inherently effective during training, eliminating the need for post-hoc pruning and preserving the model’s learned retrieval capabilities. The kernel implementation, illustrated in Figure 2 (left), is designed for efficiency. In an outer loop, the sampling weight $\Delta$ and gating coefficient $A$ are loaded into high-speed SRAM. In an inner loop, a zero-order hold method iteratively processes blocks of the value matrix $V$ from SRAM to generate content-aware masks for the key matrix $K$. These masks are then causally broadcast to match the length of the query matrix $Q$ in HBM, avoiding memory usage with quadratic complexity. Finally, the mask blocks are concatenated to form the complete content-aware dynamic mask.
4.2 Compute Position-Aware Sparse Weights
The second core component of DMA is the position-aware sparse weight computation. As outlined in Algorithm 1, this component leverages the previously generated dynamic mask to sparsify the computation of scaled dot-product attention weights, effectively reducing the per-step complexity from $O(nd_{h})$ to $O(wd_{h})$.
@@ 4.2 Compute Position-Aware Sparse Weights
For the query at step $t$, $q_{t}\in\mathbb{R}^{n_{h}\times d_{h}}$, and the complete key-value pairs, $K,V\in\mathbb{R}^{n_{h}\times n\times d_{h}}$, the entire computation flow is detailed in Equation (LABEL:eq:attention_output). Initially, for each attention head $n_{h}$, the scaled dot-product between the query and keys, $q_{t}K^{\top}$, is computed and then element-wise multiplied by the previously constructed dynamic mask $m_{t}$. The scaling factor $\sqrt{d_{h}}$ is crucial here as it prevents the dot products from becoming excessively large, which could push the softmax function into a saturated region with minimal gradients. After applying the mask, the softmax function normalizes the results to produce attention weights $p_{n_{h},j}$. Notably, when a mask value $m_{n_{h},j}=$-$\infty$, the corresponding attention weight $p_{n_{h},j}\approx 0$, effectively skipping computations for masked positions and filling them with zeros. This ensures that the model focuses solely on relevant unmasked contexts. The attention weights for each head are then multiplied by the value vectors and summed to produce the final context vector $o_{t}\in\mathbb{R}^{n_{h}\times d_{h}}$, where each row captures different contextual patterns and dependencies. The multi-head mechanism, combined with dynamic masking, allows the model to attend to various patterns in parallel across the sequence. This output integrates information from all attention heads, forming a rich hierarchical context representation that effectively captures dependencies at varying distances within the sequence history. It is important to note that this method can approximate full attention when $n_{h}\times w\leq n$, while maintaining computational efficiency.
@@
$o_{t}$ $=\operatorname{softmax}(\frac{q_{t}k^{\top}}{\sqrt{d_{h}}}\circ m_{t})v\quad where\quad p_{t}\in\mathbb{R}^{n_{h}\times n}\quad o_{t}\in\mathbb{R}^{n_{h}\times d_{h}}$ (5)